OpenStudy (anonymous):

SR12 * SR 3xy^2 Multiply the radical show work please. Im just checking my anwsers.

OpenStudy (anonymous):

6*sr(x)*y

OpenStudy (anonymous):

what about ^2?

OpenStudy (anonymous):

y^2 rite

OpenStudy (anonymous):

your confusing me....

OpenStudy (anonymous):

sr(3*x*(y)^2) rite

OpenStudy (anonymous):

that cant be the anwser....

OpenStudy (anonymous):

i was asking you the second part of que

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

2nd part is minimized to sr(3)*sr(x)*y 1st part 2*sr(3) multiply 6*sr(x)*y

OpenStudy (anonymous):

OKay then what about this one (4sr3) (-5sr21)

OpenStudy (anonymous):

-60*sr(7)

OpenStudy (anonymous):

how u get 60?

OpenStudy (anonymous):

4*(-5)*sr(3)*sr(7*3) -20*3*sr(7)

OpenStudy (anonymous):

okay this one? 5SR45x^3y^2 * sr 5x

OpenStudy (anonymous):

75*x^2*y im done

OpenStudy (anonymous):

hmm got the things huh abt sr problems

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

understood solving sr problems still any difficulty

OpenStudy (anonymous):

i understand how to do them was just checking my anwser but the ones i have trouble with are adding them. SR 24 + SR 54. do you find common bases?

OpenStudy (anonymous):

5*sr(6)

OpenStudy (anonymous):

how you get that

OpenStudy (anonymous):

24=6*4 54=6*9 sr(4)=2 sr(9)=3 so 2*sr(6)+3*sr(6) get it

OpenStudy (anonymous):

ya